Required information

[The following information applies to the questions displayed below.]

Seiko’s current salary is $133,500. Her marginal tax rate is 32 percent and she fancies European sports cars. She purchases a new auto each year. Seiko is currently a manager for Idaho Office Supply. Her friend, knowing of her interest in sports cars, tells her about a manager position at the local BMW and Porsche dealer. The new position pays only $123,000 per year, but it allows employees to purchase one new car per year at a discount of $22,900. This discount qualifies as a nontaxable fringe benefit. In an effort to keep Seiko as an employee, Idaho Office Supply offers her a $19,700 raise. Answer the following questions about this analysis.

a. What is the annual after-tax cost to Idaho Office Supply if it provides Seiko with the $19,700 increase in salary?

b-1. Financially, which offer is better for Seiko on an after-tax basis?

  • Car dealer's offer

  • Current employer's offer

  • Both offers



b-2. By how much is the offer better for Seiko on an after tax basis (Assume that Seiko is going to purchase the new car whether she switches jobs or not.)

c. What before-tax salary would Seiko need to receive from Idaho Office Supply to make her financially indifferent (after taxes) between receiving additional salary from Idaho Office Supply and accepting a position at the auto dealership? (Round your intermediate computations to the nearest dollar amount.)
Answer #1
a After tax cost
$19700-$19700*32%
$19700-$6304
$13396
b-1 Current employer's offer
b-2 current employer package
$153200 * (1-.32) = $104176
Car dealer package
$123000 * (1-.32) = $83640
$83640+$22900 = $106540
$106540-$104176 = $2364
Offer is better by $2364
c Salary received = $156676
$2364/(1-.32) = $3476
$133500+$19700+$3476 = $156676
$156676(1-.32) = $106540